artlover03:

Find the circumference and the area of a circle with radius 7ft (use pi-3.14) Circumference: Area:

michaellopez:

hold on im solving it

daddysbabygirl:

do you know the formula

michaellopez:

for circumference, use C=2πr for area, use A=πr^2

michaellopez:

do you need any more help

michaellopez:

i have the answer

daddysbabygirl:

C=2πr=2·π·7≈43.9823ft

artlover03:

thats what i got for circumference, but what about area?

daddysbabygirl:

A=πr2=π·72≈153.93804ft²

snowflake0531:

it's 14 pi.., since it's 2pir

artlover03:

is it 153.9380... or 14????

daddysbabygirl:

look at messages

snowflake0531:

14 pi

snowflake0531:

which is about 44